# Why Isn’t Relational Learning Taking Over the World?

### v1

@@ Abstract
AI seems to be taking over the world with systems that model pixels, words, and phonemes. The world is arguably made up, not of pixels, words, and phonemes but of entities (objects, things, including events) with properties and relations among them. Surely we should model these, not the perception or description of them. You might suspect that concentrating on modeling words and pixels is because all of the (valuable) data in the world is in terms of text and images. If you look into almost any company you will find their most valuable data is in spreadsheets, databases and other relational formats. These are not the form that are studied in introductory machine learning, but are full of product numbers, student numbers, transaction numbers and other identifiers that can’t be interpreted naively as numbers. The field that studies this sort of data has various names including relational learning, statistical relational AI, and many others. This paper explains why relational learning is not taking over the world – except in a few cases with restricted relations – and what needs to be done to bring it to it’s rightful prominence.
@@ Introduction
In my AI courses I often include the clicker question of Figure 1. Reading machine learning books and papers would lead one to assume the answer is A. Reading the press on generative AI would lead someone to think the answer is B. Very few of the students choose A or B.
@@
By relational learning, I mean learning models that make probabilistic predictions about entities (things, objects, including events), their properties and relations among them. The aim is to model things rather than modelling their manifestations in language or images.
Contemporary relational learning tools are typically trained to predict probabilities, often evaluated by log loss. Evaluation is much more diverse; see below.
Part of the theme of this paper is that we should not lose connection to a downstream task that the learned model is used for. Learning is usually not an end in itself, but is used for decision making. This could be for autonomous decision making, where probabilities and utilities are needed, or human-supported decision making where multiple scenarios may be useful as long as they can be evaluated by a human decision maker.
@@
If you were to examine company data, environmental data, or government data – with few exceptions – you will find that the most valuable data not in natural language text or images but in spreadsheets, relational databases and other relational formats. This relational data is typically not the sort of tabular data used in introductory machine learning courses – for which standard gradient tree boosting such as XGBoost (Chen and Guestrin 2016) works directly on – but is full of product numbers, student numbers, transaction numbers and other identifiers. These identifiers are often represented as numbers, but the numbers themselves don’t convey meaning. Gradient tree boosting can be effective when made aware of relational structure (Natarajan et al. 2012). Identifiers are a standard tool in the database design toolkit.
@@
  • Does the system know what exists? Most databases assume that they know what exists; things that exist are given names or identifiers, and what doesn’t exist is ignored. In many real-world cases, we don’t know what exists. What there a person in the house last night? Who threw the rock that broke the window? (It might have not been a rock that broke the window, and maybe no one threw it.)
@@
The second way to converting relations to triples is to use (row, column, value) triples. The row can be a primary key if there is one, or is a new identifier – the “reified” entity. Reify means “make into an entity” (or “thingify”). This is done in Figure 2 (b), where q262802-3B1AE42E and Q262802-9c5c267f are reified entities. The number of facts about a reified entity is number of columns in table, which is typically just a handful. To convert to the original relation, one can do a database join of the triples on the reified entities and project onto the other properties (which is how I generated the table from the Wikidata triples). Reification does not lose information, but make the representation more modular: other properties, including meta-properties (such as the author) can be easily added to a knowledge graph.
Another way to convert relations into triples is to user pairs of the attributes in a star-to-clique method (Wen et al. 2016). This, however, loses information.
@@
We [^2] Bahare Fatemi, Mehran Kazemi, Ali Mehr and Martin Wang. See https://cs.ubc.ca/~poole/res/FB15k-triples.txt for example triples for all 1344 predicates. created English version of the test set that, together with the plausible examples that existing methods ranked highly, could be used to crowdsource ground truth. The ground truth could then be used to evaluate prediction methods. However, we decided determining whether a tuple was true wasn’t a skill people could do quickly and accurately, even when they had access to Wikipedia, Google and other online sources. Positive examples were straightforward, but determining something is false requires guessing.
Toutanova and Chen (2015) created a variant of FB15k called FB15k-237. They noticed that a predictor could do reasonably well on FB15k by just determining that two predicates were approximately equivalent or inverses. To determine whether a predictor was doing more than this, they created FB15k-237 with the inverse functions removed (it would remove one first two triples of Figure 3). FB15k-237 isn’t better than FB15k; they test different things. Together they provide interesting evidence: the difference between them tell us whether a predictor can recognize inverses, and the performance on FB15k-237 tells us whether the method can exploit other regularities. We would want a predictor that can determine inverses, but would like it to do more than that.
@@
Big data often implies more entities but less data, on average, about each entity. This happens when when more detail is added about each entity as well as when more entities are added. As the analysis in Appendix A shows, the top-level entities about which there is more data tend to have more associated reified entities (about which there is very little data). There is invariably also a long tail of entities about which not much is known.
@@
There are similar problems with standard datasets in the graph learning community (Bechler-Speicher et al. 2025; Roy et al. 2025). JF17K (Wen et al. 2016),
FB-AUTO, M-FB15K (Fatemi et al. 2020) and RelBench (Robinson et al. 2024) are benchmarks that do not convert to triples. Hopefully such datasets can replace the older benchmarks, but still more needs to be done.
@@
Surprisingly, there is very little work on making predictions about the future from arbitrary complete databases. We can’t add an airline reservation system to a learner and get it to predict future demand. It might be because most company databases are proprietary and include much metadata, such as the database schema. Individual companies such as Amazon and Netflix have huge datasets they are mining, but the techniques are proprietary and are probably tuned to the particular task, such as maximizing engagement, rather than allowing general queries of what is predicted. Steck et al. (2021) describe techniques used by Netflix. Feature engineering (such as carefully chosen joins and projections) can created tables that methods such as XGBoost work on.
@@
Databases do not store arbitrary relations. Some relations are stored and others are derived. Which relations to store is carefully designed (Codd 1970). The lack of negative statement is part of database design. Database designers choose to store only the true statements for some of the possible relations, selected to make the database small. They don’t represent “did-not-buy” or “has-watched-movie-starring” as explicit relations, instead choose smaller relations to explicitly represent, and use database queries and rules to infer other ones.
@@
  • parameter sharing or weight typing; namely “logical variables” and “quantification” in logic approaches (Poole 1993; De Raedt, Kimmig, and Toivonen 2007; Domingos and Lowd 2009), exchangeability [^3] Identifiers can be exchanged and the probability distribution should not change. Exchangeability of identifiers does not imply exchangeability of random variables (de Finetti 1974). (Niepert and Van den Broeck 2014), or the “convolutional” in convolutional graph neural networks (Schlichtkrull et al. 2018).
  • aggregation, where a property or an entity or relation among entities depends on other entities, such as predicting the gender or age of a person from the movies they have watched.
@@
The evaluation of a knowledge graphs depends on what is asked. To predict whether a random triple is in Wikidata, answering “no” is over 99.99999999995% accurate [^4] Wikidata has about 1.65 billion triples on 117 million items and 290,000 properties (https://grafana.wikimedia.org/d/000000175/wikidata-datamodel-statements). With no domain or range constraints in Wikidata, any entity can be the subject or object of any property. (Let’s ignore strings and other datatypes as values, as there are too many possible values). There are thus $(1.17e8)^{2}*2.9e5\approx 4e21$ possible triples, of which $131.6e9/4e21=4e{-}$ 13 are in Wikidata. The number of true triples is difficult to estimate; if Wikidata contains 0.01% of the true triples, predicting “no” would be over 99. 9999995% accurate for predicting truth.. Accuracy of random triples is not a reasonable measure. Most random triples are nonsense, which no one would ask and could be answered just by knowing the types of the entities.
@@
The most common way to evaluate a knowledge graph predictor is to use ranking: given test triple $(s,v,o)$, ask the predictor to rank the entities related to a given subject $s$ and verb $v$, i.e., to predict which triples of the form $(s,v,?)$ are true (and similarly to rank the subjects for a given verb and object). Given a total ordering of predictions $o_{1},o_{2},o_{3},\dots$, the rank is the $i$ such that $o=o_{i}$. In filtered prediction (Bordes et al. 2013b), the $o_{i}$ that are known to be true (because the corresponding triple is in the training or test set) are removed from the ordering.
The mean rank is one possible measure but is dominated by a few very large ranks, and is not used much any more. A common measure is “hit-at-$k$” for some integer $k$ (such as 1, 3 or 10), the proportion of test cases where the rank of the correct answer is $k$ or less. The $k$ in hit-at-$k$ seems like an arbitrary parameter. A method that is common and seems to not have arbitrary parameters is the mean reciprocal rank (MRR) (Radev et al. 2002). This weights the $k$th rank as $1/k$. It can be interpreted as an agent who always considers the first item, considers the second item half the time, the 3rd item 1/3 of the time on so on. However this interpretation is problematic; see Appendix B.
@@
  • The actual probabilities are lost. One predictor might be very sure of one answer, and have to fill in the rest. Another predictor might have a handful of plausible answers, they are not sure about. These predictors could produce the same ordering. There is no reward for being sure (and correct) or for actually specifying uncertainty, even though this is valuable information for making decisions.
  • It loses sight of the downstream task. Considering top-$k$ is reasonable for cases where the results can be checked, such as in movie recommendation, web search or suggesting to a doctor the $k$ most likely diagnoses so they can be reminded of something they may have overlooked. It seems less useful when just presented $k$ possible answers for which there is no easy way to check which, if any, are correct. State-of-the-art methods that combines existing models (Li et al. 2024; Jeon, Lim, and Suk Choi 2024) have hit-at-10 rates on FB15k-237 of 55.8%. and 52.58% respectively, which does not seem very useful for any real-world task where correctness cannot be easily verified.
@@
There are many success stories from instances of relational learning from protein prediction (Jumper et al. 2021), where the atoms are entities and relations includes bonds, to predicting congestion on roads and at intersections for route-planning (Delling et al. 2015) to the combination of ML and the internet of things in agriculture (Sharma and Shivandu 2024; Mureithi 2024). These build models particular to the application. Generalizing particular models so we can make predictions from arbitrary relational datasets and queries is still in its infancy, despite the years of research. This provides a great opportunity for researchers to make fundamental contributions. For relational learning to reach its full potential, the following (at least) need to be done:
@@
For incomplete knowledge bases, we need to explicitly consider why data is missing. Having test examples that are a random subset of all of the tuples, does not provide a surrogate to real-world problems where data is invariably not missing at random (Rubin 1976). I doubt there are any Taylor Swift officially released albums missing from Wikidata, but for most of the recording artists on Wikidata most of their recordings are missing. Probabilistic models of missing data (Marlin et al. 2011; Mohan, Pearl, and Tian 2013) tend to be heavyweight, requiring detailed model of why data is missing. This is difficult when virtually all of the tuples are missing. We made an attempt to build a lightweight model of missing data (Poole, Mehr, and Wang 2020) with the aim of extended it to relational models.
@@
To realize the full potential of relational learning, we need to combine the information from all datasets, and build predictive models from these. Combining multiple heterogenous datasets is the challenge of the semantic web (Berners-Lee, Hendler, and Lassila 2001). Building predictive models from heterogenous data might seem impossible, but we know how to; it’s called science. To combine all recorded experimental and observational data to make better predictions, we need at least ontologies [^6] Scientific ontologies include Open Biological and Biomedical Ontology http://obofoundry.org, WHO Family of International Classifications https://www.who.int/standards/classifications, SMOMED-CT healthcare ontology https://www.snomed.org (to allow semantic interoperability) and provenance information (Gil et al. 2016; Sikos, Seneviratne, and McGuinness 2021), including what was manipulated, what was controlled for, what was observed and when, how was what to record decided (modelling the missing information), who manipulated the data, etc. There needs to be multiple hypotheses, including the null hypothesis (the data is noise) and that some data is erroneous or entirely fictitious (to counter misinformation). Multiple hypotheses may need to be combined to make a prediction for an actual case (Poole 2012). With an interconnected world, we don’t need to make conclusions from individual experiments, but can learn from everyones experiences, which should enable more accurate hypotheses.
@@ Appendix A Appendix A: Proportion of Wikidata entities with few tuples
Wikidata ID # tuples #$\#<10$ name

### v2

@@ Abstract
Artificial intelligence seems to be taking over the world with systems that model pixels, words, and phonemes. The world is arguably made up, not of pixels, words, and phonemes but of entities (objects, things, including events) with properties and relations among them. Surely we should model these, not the perception or description of them. You might suspect that concentrating on modeling words and pixels is because all of the (valuable) data in the world is in terms of text and images. If you look into almost any company you will find their most valuable data is in spreadsheets, databases and other relational formats. These are not the form that are studied in introductory machine learning, but are full of product numbers, student numbers, transaction numbers and other identifiers that can’t be interpreted naively as numbers. The field that studies this sort of data has various names including relational learning, statistical relational AI, and many others. This paper explains why relational learning is not taking over the world – except in a few cases with restricted relations – and what needs to be done to bring it to it’s rightful prominence.
@@ Introduction
In my AI courses I often include the clicker question of Figure 1. Reading machine learning books and papers would lead one to assume the answer is A. Reading the press on generative AI would lead one to think the answer is B. Very few of the students choose A or B.
@@
By relational learning, I mean learning models that make probabilistic predictions about entities (things, objects, including events), their properties and relations among them. The aim is to model entities rather than modelling their manifestations in language or images.
Contemporary relational learning tools are typically trained to predict probabilities, optimizing log loss. Evaluation is much more diverse; see below.
Part of the theme of this paper is that we should not lose connection to the downstream task that a learned model is used for. Learning is usually not an end in itself, but is used for decision making. This could be for autonomous decision making, where probabilities and utilities are needed, or human-supported decision making where multiple scenarios can be evaluated by a human decision maker.
@@
If you were to examine company data, environmental data, or government data – with few exceptions – you will find that the most valuable data not in natural language text or images but in spreadsheets, relational databases and other relational formats. This relational data is typically not the sort of tabular data used in introductory machine learning courses – for which standard algorithms such as gradient tree boosting (Chen and Guestrin 2016) work directly on – but is full of product numbers, student numbers, transaction numbers and other identifiers. These identifiers are often represented as numbers, but the numbers themselves don’t convey meaning. For gradient tree boosting to be effective for such data, it needs to exploit relational structure (Natarajan et al. 2012).
@@
  • Does the system know what exists? Most databases assume that they know what exists; things that exist are given identifiers, and what doesn’t exist is ignored. In many real-world cases, we don’t know what exists. What there a person in the house last night? Who threw the rock that broke the window? (It might have not been a rock or maybe no one threw it.)
@@
The second way to converting relations to triples is to use (row, column, value) triples. The row can be a primary key if there is one, or is a new identifier – the “reified” entity. Reify means “make into an entity” (or “thingify”). In Figure 2 (b), Q262802-3B1AE42E and Q262802-9c5c267f are reified entities. The number of facts about a reified entity is the number of columns in the table, which is typically just a handful. To convert to the original relation, one can do a database join of the triples on the reified entities and project onto the other properties (which is how I generated the table from the Wikidata triples). Reification does not lose information, but make the representation more modular: other properties, including meta-properties (such as the author) can be easily added to a knowledge graph.
Another way to convert relations into triples is to pair attributes, called the star-to-clique method (Wen et al. 2016). This, however, loses information. Figure 3 shows some examples of the sort of triples that arise from this method.
@@
We [^2] Bahare Fatemi, Mehran Kazemi, Ali Mehr and Martin Wang. See https://cs.ubc.ca/˜poole/res/FB15k-triples.txt for example triples for all 1344 predicates. created English version of the test set that, together with the plausible examples that existing methods ranked highly, could be used to crowdsource ground truth. The ground truth could then be used to evaluate prediction methods. However, we decided determining whether a tuple was true wasn’t a skill people could do quickly and accurately, even when they had access to Wikipedia, Google and other online sources. Positive examples were straightforward, but determining something is false requires guessing.
Toutanova and Chen (2015) noticed that a predictor could do reasonably well on FB15k by just determining that two predicates were (approximately) equivalent or inverses. To determine whether a predictor was doing more than this, they created FB15k-237 with the inverse functions removed (it would remove one first two triples of Figure 3). FB15k-237 isn’t better than FB15k; they test different things. Together they provide interesting evidence: the difference tells us whether a predictor can recognize inverses; the performance on FB15k-237 tells us whether the method can exploit other regularities. We would want a predictor that can determine inverses, but would like it to do more than that.
@@
Big data often implies more entities, but less data, on average, about each entity. As entities are added, they start off as stubs with few tuples, resulting in a long tail of entities about which not much is known. As more detail is added about an entity, there are more associated reified entities, about which there are inherently few tuples. See Appendix A.
@@
Similar problems with standard benchmarks arise in the graph learning community (Bechler-Speicher et al. 2025; Roy et al. 2025). Although relational learning can be seen as an instance of graph learning, and graph learning an instance of relational learning, the challenges in this paper do not seem to have much in common with generic graph tasks.
JF17K (Wen et al. 2016), FB-AUTO, M-FB15K (Fatemi et al. 2020) and RelBench (Robinson et al. 2024) are benchmarks that do not convert to triples. Such datasets are a start in building more realistic test sets.
@@
Surprisingly, there is very little work on making predictions about the future from arbitrary complete databases. We can’t add an airline reservation system to a learner and get it to predict future demand. It might be because most company databases are proprietary and include much metadata, such as the database schema. Individual companies such as Amazon and Netflix have huge datasets they are mining, but the techniques are proprietary and are undoubtedly tuned to the particular task, such as maximizing profit and engagement, rather than allowing general queries of what is predicted. Steck et al. (2021) describe techniques used by Netflix. Feature engineering (such as carefully chosen joins and projections) can created tables that methods such as XGBoost work on.
@@
Databases do not store arbitrary relations. Some relations are stored and others are derived. Which relations to store is carefully designed (Codd 1970). Database designers choose to store only true statements for some of the possible relations, selected to make the database small. They don’t represent “did-not-buy” or “has-watched-movie-starring” as explicit relations, instead choose smaller relations to explicitly represent, and use queries and rules to infer other ones.
@@
  • parameter sharing or weight tying; namely “logical variables” and “quantification” in logic approaches (Poole 1993; De Raedt, Kimmig, and Toivonen 2007; Domingos and Lowd 2009), exchangeability [^3] Identifiers can be exchanged and the probability distribution should not change. Exchangeability of identifiers does not imply exchangeability of random variables (de Finetti 1974). (Niepert and Van den Broeck 2014), or the “convolutional” in convolutional graph neural networks (Schlichtkrull et al. 2018).
  • aggregation, where a property of an entity or relation among entities depends on other entities, such as predicting the gender or age of a person from the movies they have watched.
@@
[…]
@@
[…]
@@
  • The actual probabilities are lost. One predictor might be very sure of one answer, and have to fill in the rest. Another predictor might have a handful of plausible answers, they are not sure about. These predictors could produce the same ordering. There is no penalty for being overconfident or reward for being sure (and correct) or for actually specifying uncertainty, even though this is valuable information for making decisions.
[…]
@@
There are many success stories from instances of relational learning from protein prediction (Jumper et al. 2021), where the atoms are entities and relations includes bonds, to predicting congestion on roads and at intersections for route-planning (Delling et al. 2015) to the combination of ML and the internet of things in agriculture (Sharma and Shivandu 2024; Mureithi 2024). These build models particular to the application. Generalizing particular models so we can make predictions from arbitrary relational datasets and queries is still in its infancy, despite years of research. This provides a great opportunity for researchers to make fundamental contributions. For relational learning to reach its full potential, the following (at least) need to be done:
@@
For incomplete knowledge bases, we need to explicitly consider why data is missing. Having test examples that are a random subset of all of the tuples, does not provide a surrogate to real-world problems where data is invariably not missing at random (Rubin 1976). I doubt there are any Taylor Swift officially released albums missing from Wikidata, but for most of the recording artists on Wikidata most of their recordings are missing. Probabilistic models of missing data (Marlin et al. 2011; Mohan, Pearl, and Tian 2013) tend to be heavyweight, requiring detailed modelling of why data is missing. This is difficult when virtually all of the tuples are missing. We made an attempt to build a lightweight model of missing data (Poole, Mehr, and Wang 2020) with the aim of extended it to relational models.
@@
To realize the full potential of relational learning, we need to combine the information from all datasets, and build predictive models from these. Combining multiple heterogenous datasets is the challenge of the semantic web (Berners-Lee, Hendler, and Lassila 2001). Building predictive models from heterogenous data might seem impossible, but we know how to; it’s called science. To combine all recorded experimental and observational data to make better predictions, we need at least ontologies [^6] Scientific ontologies include Open Biological and Biomedical Ontology http://obofoundry.org, WHO Family of International Classifications https://www.who.int/standards/classifications, SMOMED-CT healthcare ontology https://www.snomed.org (to allow semantic interoperability) and provenance information (Gil et al. 2016; Sikos, Seneviratne, and McGuinness 2021), including what was manipulated, what was controlled for, what was observed and when, how was what to record decided (modelling the missing information), who manipulated the data, etc. Science involved building hypotheses, typically from multiple datasets, that persist, ready to be revised or rejected as new evidence comes in. Hypotheses can be about anything; from the effects of global warming, to what foods are edible, to the traffic in Vancouver. There needs to be multiple hypotheses for a dataset, including the null hypothesis (the data is noise) and hypotheses that some data is erroneous or entirely fictitious (to counter misinformation). Multiple hypotheses may need to be combined to make a prediction for an actual case (Poole 2012). With an interconnected world, we don’t need to make conclusions from individual experiments, but can learn from everyones experiences, which should enable more accurate hypotheses.
@@ Appendix A Appendix A: Proportion of Wikidata entities with few tuples
Wikidata ID # tuples #<10\#<10# < 10 name